# Efficient Model-Based Reinforcement Learning Through Optimistic Thompson Sampling

### v1

@@ Appendix C Evaluation Tasks
$r(s,a)=-\rho\left(\exp\left(-\sum_{i=1}^{6}a_{i}^{2}\right)-1\right)+\begin{cases}\exp\left(-\left\|\text{EE}(s)-\text{goal}\right\|^{2}\right)&\text{if }\left\|\text{EE}(s)-\text{goal}\right\|<\epsilon\\ 0&\text{otherwise}\end{cases}.$
@@ Appendix D Extended Experiments
In this section, we extend our experiments by evaluating H-UCRL approx and MBPO without uncertainty clamping on the Half-Cheetah task, with the results given in Figure 4.
@@
As described in Section A, H-UCRL approx differs from the algorithm presented in Curi et al. (2020) in a few ways: (1) we use our GP model rather than a probabilistic ensemble, (2) we use the sampling-based optimistic hallucination method in equation 6 to select states, and (3) we use predicted rewards to guide optimism rather than the ground truth rewards. This version of H-UCRL does not perform well on Half-Cheetah relative to MBPO and HOT-GP, with the best seeds achieving mean episode rewards as high as 6000 after training with 250,000 environment steps. As we observed when discussing H-UCRL performance on the coverage task, this performance gap may stem from infeasible optimism, where the rewards associated with perturbed states $\mu(s,a)+\beta\ \Sigma(s,a)\eta$ are high but unreachable in practice. Additionally, a suboptimal choice of $\beta$ could be a contributing factor, as we did not fine tune this parameter on Half-Cheetah due to its extended runtime and instead used the optimal $\beta$ finetuned on the coverage task. HOT-GP also has a hyperparameter $r_{\text{min}}$, which controls the level of optimism during training. While the domain dependence of $r_{\text{min}}$ means some adaptation is required for new tasks, we found that setting $r_{\text{min}}$ is generally intuitive and, moreover, provides interesting insights into the nature of the task itself. For instance, if the dynamics in the domain are simple, we can be more optimistic about exploration (i.e., use a higher value of $r_{\text{min}}$), whereas if they are complex, we need to spend more time learning the dynamics before becoming more optimistic.

### v2

@@ Appendix C Evaluation Tasks
$r(s,a)=-\rho\left(\exp\left(-\sum_{i=1}^{6}a_{i}^{2}\right)-1\right)+\begin{cases}\exp\left(-\left\|\text{EE}(s)-\text{goal}\right\|^{2}\right)&\text{if }\left\|\text{EE}(s)-\text{goal}\right\|<\epsilon\\ 0&\text{otherwise.}\end{cases}$
@@
Figure 5: Visualizations of the U Maze (left) and Medium Maze (right) environments with randomly selected goal locations.
U Maze. In this task, the agent must navigate through a U-shaped maze shown in Figure 5 to reach a randomized target location. The agent receives a sparse reward of 1 when it comes within a small threshold distance $\epsilon$ of the target and 0 otherwise. Specifically, the reward function is
$r(s,a)=\begin{cases}1&\text{if $||\text{Pos}(s)-\text{goal}||<\epsilon$}\\ 0&\text{otherwise.}\end{cases}$
The agent observes its position $(x,y)$ and 2-dimensional goal position. Actions are 2-dimensional vectors representing velocity in the $x$ and $y$ dimensions.
Medium Maze. This task shares the same observation and reward structure with the U Maze. The only difference is that this maze environment is significantly more complex, as shown in Figure 5.
@@ Appendix D Extended Experiments
In this section, we extend our experiments in order to better contextualize HOT-GP. First, we investigate the asymptotic performance of HOT-GP and all comparison methods on the Sparse Reacher task. We also provide learning curves on the Sparse Reacher task with action penalty 0.3 to supplement the results in Figure 1) with an intermediate action penalty. Finally, we evaluate the approximation of H-UCRL and MBPO without uncertainty clamping on the Half-Cheetah task.
Sparse Reacher. To supplement the Sparse Reacher results presented in Figure 1, we provide an additional learning curve on the Sparse Reacher using action penalty of 0.3. These results are consistent with the trends observed using other action penalties and HOT-GP variants perform the best.
Figure 6: Learning curves on the Sparse Reacher with action penalty of 0.3. HOT-GP achieves superior sample efficiency and performance. The dashed line denotes SAC performance at convergence within 1,000,000 environment steps.
@@
Half-Cheetah. As described in Section A, H-UCRL approx differs from the algorithm presented in Curi et al. (2020) in a few ways: (1) we use our GP model rather than a probabilistic ensemble, (2) we use the sampling-based optimistic hallucination method in equation 6 to select states, and (3) we use predicted rewards to guide optimism rather than the ground truth rewards. This version of H-UCRL does not perform well on Half-Cheetah relative to MBPO and HOT-GP, with the best seeds achieving mean evaluation episode rewards as high as 6000 after training with 250,000 environment steps. As we observed when discussing H-UCRL performance on the coverage task, this performance gap may stem from infeasible optimism, where the rewards associated with perturbed states $\mu(s,a)+\beta\ \Sigma(s,a)\eta$ are high but unreachable in practice. Additionally, a suboptimal choice of $\beta$ could be a contributing factor, as we did not fine tune this parameter on Half-Cheetah due to its extended runtime and instead used the optimal $\beta$ finetuned on the coverage task. HOT-GP also has a hyperparameter $r_{\text{min}}$, which controls the level of optimism during training. While the domain dependence of $r_{\text{min}}$ means some adaptation is required for new tasks, we found that setting $r_{\text{min}}$ is generally intuitive and, moreover, provides interesting insights into the nature of the task itself. For instance, if the dynamics in the domain are simple, we can be more optimistic about exploration (i.e., use a higher value of $r_{\text{min}}$), whereas if they are complex, we need to spend more time learning the dynamics before becoming more optimistic.